OpenStudy (anonymous):

X^3+512=0 real or imaginary solution by factoring

OpenStudy (jdoe0001):

\(\bf \Large x^3+512=0\implies x^3=-512\implies x = \sqrt[3]{-512}\) what do you think?

jimthompson5910 (jim_thompson5910):

it's better to use the sum of cubes factoring formula (so you can find the imaginary solutions as well)

OpenStudy (jdoe0001):

hmm ohh yeah

jimthompson5910 (jim_thompson5910):

x^3+512 = 0 x^3 + 8^3 = 0 (x+8)(x^2 - x*8 + 8^2) = 0 ... use the sum of cubes factoring formula (x+8)(x^2 - 8x + 64) = 0 I'll let you finish up